# Head-Tail-Aware KL Divergence in Knowledge Distillation for Spiking Neural Networks

### v1

@@ Abstract
Spiking Neural Networks (SNNs), regarded as the third generation of neural networks, offer a biologically inspired framework for information processing in neural networks. They transmit information via discrete spikes, enabling inherently low-latency and energy-efficient computation. This makes SNNs particularly well-suited for deployment on neuromorphic computing hardware platforms [1, 2, 3, 4]. Despite their biological plausibility and energy efficiency, SNNs currently underperform compared to traditional Artificial Neural Networks (ANNs) in tasks such as image classification, object segmentation, and natural language processing. Due to the non-continuous and non-differentiable nature of spike-based activation functions, SNNs cannot be directly trained using the standard Backpropagation (BP) algorithm widely used in ANNs. Instead, they rely on alternative training techniques such as Spatio-Temporal Backpropagation (STBP) or Surrogate Gradient(SG)[5, 6].
@@ II-A SNN Training Methods
SNN training is primarily categorized into two approaches: ANN-to-SNN conversion and direct training of SNNs. The conversion-based approach transfers knowledge from pre-trained ANNs to SNNs by substituting ReLU activations with spike-based neuronal models. However, such methods often fall short in fully leveraging the spatio-temporal dynamics inherent to SNNs, limiting their potential in tasks where temporal encoding is critical[13, 14]. Direct SNN training, which uses spike-based temporal information, avoids these limitations[15]. Innovations like tdBN[16] and MPBN[17] improve training by optimizing feature normalization and stabilizing membrane potential updates. ESL-SNNs[18] propose an efficient evolutionary structure learning framework for SNNs. Several modifications to ResNet architectures, including SEW-ResNet[19], DS-ResNet[20], and MS-ResNet [14], have also been explored to enhance performance. Additionally, networks like RSNN[21] enhances spatiotemporal processing by using a recurrent structure to preprocess input slices. The integration of Transformer-based SNNs is another area of focus, with models like[22, 23, 24] improving classification accuracy using spike-based self-attention mechanisms. STMixer[25] models is designed for event-driven asynchronous chips.
@@ II-B Knowledge Distillation for SNNs
KD is a widely adopted model compression method, where a smaller student model is trained to replicate the behavior of a larger, more accurate teacher model. This strategy allows the student model to maintain competitive accuracy while significantly reducing computational cost and model size[26, 27, 28]. In the context of SNNs, KD has emerged as a critical approach for enhancing performance[29]. Typically, a pre-trained ANN acts as the teacher, offering robust feature extraction capabilities, while the SNN serves as the student, preserving its inherent advantages such as event-driven computation and low energy consumption. This methodology facilitates knowledge transfer from ANN to SNN, effectively combining their complementary strengths.
Recent advancements have significantly expanded the scope and effectiveness of KD in SNNs. Distilling Spikes[30] proposes a multi-stage distillation framework employing intermediate networks to improve student SNN performance. Spike-Thrift[31] reduces spike activity via attention-guided ANN compression and sparse training. LaSNN[10] introduces a layer-wise distillation framework incorporating attention mechanisms to bridge the representation and transmission gap between ANNs and SNNs. KDSNN[9] presents a unified distillation strategy that combines both logits-based and feature-based knowledge transfer. SpikeBERT[32] adopts a two-phase distillation approach to transfer knowledge from BERT into a spiking language model. Joint A-SNN[33] integrates self-distillation and weight factorization to further enhance model performance. [34] proposed an evolutionary method that combines knowledge distillation and pruning to optimize SNN structures. The SAKD[35] applies a dual-phase training process, weight transfer and behavioral imitation. BKDSNN[11] framework enhances feature transferability through blurred knowledge distillation, yielding improved performance on complex datasets. Despite these advancements, fully leveraging the spatio-temporal dynamics of spikes remains a significant challenge. Thus, the design of distillation strategies explicitly tailored for the unique characteristics of SNNs remains a key direction for future research, with the potential to achieve further improvements in both accuracy and energy efficiency.

### v2

@@ Abstract
Spiking Neural Networks (SNNs), regarded as the third generation of neural networks, offer a biologically inspired framework for information processing in neural networks. They transmit information via discrete spikes, enabling inherently low-latency and energy-efficient computation. This makes SNNs particularly well-suited for deployment on neuromorphic computing hardware platforms[1, 2, 3, 4, 5, 6, 7]. Despite their biological plausibility and energy efficiency, SNNs currently underperform compared to traditional Artificial Neural Networks (ANNs) in tasks such as image classification, object segmentation, and natural language processing. Due to the non-continuous and non-differentiable nature of spike-based activation functions, SNNs cannot be directly trained using the standard Backpropagation (BP) algorithm widely used in ANNs. Instead, they rely on alternative training techniques such as Spatio-Temporal Backpropagation (STBP) or Surrogate Gradient(SG)[8, 9].
@@ II-A SNN Training Methods
SNN training is primarily categorized into two approaches: ANN-to-SNN conversion and direct training of SNNs. The conversion-based approach transfers knowledge from pre-trained ANNs to SNNs by substituting ReLU activations with spike-based neuronal models. However, such methods often fall short in fully leveraging the spatio-temporal dynamics inherent to SNNs, limiting their potential in tasks where temporal encoding is critical[16, 17]. Direct SNN training, which uses spike-based temporal information, avoids these limitations[18]. Innovations like tdBN[19] and MPBN[20] improve training by optimizing feature normalization and stabilizing membrane potential updates. Frequency-based attention mechanisms have been introduced to optimize intermediate spike features and reduce redundant spiking[21]. ESL-SNNs[22] propose an efficient evolutionary structure learning framework for SNNs. Several modifications to ResNet architectures, including SEW-ResNet[23], DS-ResNet[24], and MS-ResNet [17], have also been explored to enhance performance. Additionally, networks like RSNN[25] enhances spatiotemporal processing by using a recurrent structure to preprocess input slices. The integration of Transformer-based SNNs is another area of focus, with models like[26, 27, 28] improving classification accuracy using spike-based self-attention mechanisms. STMixer[29] models is designed for event-driven asynchronous chips.
@@ II-B Knowledge Distillation for SNNs
KD is a widely adopted model compression method, where a smaller student model is trained to replicate the behavior of a larger teacher model. This strategy allows the student model to maintain competitive accuracy while significantly reducing computational cost and model size[30, 31, 32]. As a critical approach for enhancing performance in SNNs, a pre-trained ANN acts as the teacher, providing robust feature extraction capabilities, while the SNN serves as the student, retaining its inherent advantages such as event-driven computation and low energy consumption[33]. Distilling Spikes[34] proposes a multi-stage distillation framework employing intermediate networks to improve student SNN performance. Spike-Thrift[35] reduces spike activity via attention-guided ANN compression and sparse training. LaSNN[13] introduces a layer-wise distillation framework incorporating attention mechanisms to bridge the representation and transmission gap between ANNs and SNNs. KDSNN[12] presents a unified distillation strategy that combines both logits-based and feature-based knowledge transfer. SpikeBERT[36] adopts a two-phase distillation approach to transfer knowledge from BERT into a spiking language model. Joint A-SNN[37] integrates self-distillation and weight factorization to further enhance model performance. [38] proposed an evolutionary method that combines knowledge distillation and pruning to optimize SNN structures. The SAKD[39] applies a dual-phase training process, weight transfer and behavioral imitation. BKDSNN[14] framework enhances feature transferability through blurred knowledge distillation, yielding improved performance on complex datasets. Despite these advancements, fully leveraging the spatio-temporal dynamics of spikes remains a significant challenge. Thus, the design of distillation strategies explicitly tailored for the unique characteristics of SNNs remains a key direction for future research, with the potential to achieve further improvements in both accuracy and energy efficiency.